what is the wavelength of the x-ray emitted from a tungsten target (Z=74) when an electron moves from the M-shell (n=3) to the K-shell(n=1)
1
Expert's answer
2020-06-09T13:19:04-0400
"\frac{1}{\lambda}=R(Z-1)^2\left(\frac{1}{1^2}-\frac{1}{3^2}\right)"

"\frac{1}{\lambda}=1.097\cdot10^7(74-1)^2\left(\frac{1}{1^2}-\frac{1}{3^2}\right)"

"\lambda=1.9\cdot10^{-11}\ m"